OpenStudy (anonymous):

Mel Doane is taking his office staff out to lunch. He has left on his credit card a spending limit of $99.50. If the total meal includes 2 fruitcups, 2 onion soups, 2 veal, 2 roast beef, 3 coffees and 1 soft drink, will Mel be able to use the card? (Disregard tip plus tax). If so, how much more charging will Mel be allowed before hitting his credit limit?

OpenStudy (anonymous):

the fruitcup is 1.85 and onion soups 2.95 and veal 10.45 and roast beef is 11.95 and the coffees is .95 and soft drink 1.50

OpenStudy (anonymous):

can someone help me

OpenStudy (anonymous):

add up and see whether it is greater or less than $99.50

OpenStudy (anonymous):

\[2\times 1.85+2\times 2.95+2\times 10.45+2\times 11.95+3\times .95+1.5\]